# Machine Learning Detection of Lithium Plating in Lithium-ion Cells: A Gaussian Process Approach

### v3

@@ V Conclusions
We have presented a voltage-only method for lithium plating detection that employs an exact Gaussian Process (GP) prior for capacity as a function of voltage, $Q(V)$. By analytically computing $dQ/dV$ from the joint value–derivative posterior, the method avoids the noise amplification inherent to finite-difference ICA/DVA and additionally provides calibrated uncertainty as a reliability metric.
Across all tested conditions, the GP-based analysis yields a clear separation between plating and non-plating cases based on the location of the dominant incremental-capacity peak. Cycles whose $dQ/dV$ maximum occurs below $4.0~\mathrm{V}$ are consistently plating-negative, whereas cycles with a dominant peak above $4.0~\mathrm{V}$ near end-of-charge show plating. Baseline conditions ($1\mathrm{C}$ at $25^{\circ}\mathrm{C}$ and $40^{\circ}\mathrm{C}$) exhibit no high-confidence peaks above $4.0~\mathrm{V}$, while colder and/or higher-rate conditions produce pronounced $4.0~\mathrm{V}$ peaks with tight uncertainty bands, consistent with plating.
These differential signatures align closely with independent performance metrics. Plating-positive cases exhibit faster declines in delivered charge throughput per CC cycle as well as accelerated capacity fade confirmed by periodic RPTs, with the coldest and highest-rate cases degrading most severely. In contrast, baseline and non-plating cases show stable charge throughput and much slower capacity fade.
Taken together, the concurrence of (i) high-confidence $dQ/dV$ peaks localized above $4.0~\mathrm{V}$, (ii) reduced charge throughput over cycling, and (iii) accelerated capacity fade provides a coherent, physics-consistent picture linking the GP-identified signature to practically relevant performance loss. These results demonstrate that exact GP–based voltage analysis enables more robust lithium plating diagnostics, paving the way for reliable online detection and integration into advanced battery management systems.
@@ ACKNOWLEDGMENT
We appreciate the funding support from the USAF Proteus Space STTR and the NSF CAREER Program (Grant No.2046292).

### v4

@@ V Conclusions
We presented a voltage-base method for lithium plating detection that employs Gaussian Process. By analytically computing $dQ/dV$ from voltage and current using the joint value–derivative posterior of GP, the method avoids the noise amplification inherent to finite-difference ICA/DVA and provides calibrated uncertainty as a reliability metric. Across the tested conditions, the GP-based analysis separates cases that exhibit a distinct high-voltage secondary peak feature from cases that do not. Feature-positive cases span a range of conditions, including room-temperature charging at 1C while the feature-negative cases demonstrate that the absence of plating depends on the specific electrochemical state rather than a simple charging rate–temperature threshold.
The GP-resolved secondary $dQ/dV$ peak above 4.0 V serves as the primary indicator of plating onset, with reduced charge throughput and accelerated capacity fade providing corroborating evidence in cases where plating has progressed. Notably, the method detects plating even in conditions such as 0.2C at $0\,^{\circ}\text{C}$ or 1C at $25\,^{\circ}\text{C}$ where degradation metrics remain near non-plating levels, demonstrating the sensitivity of the GP-based approach as an early-warning diagnostic. Moreover, the method demonstrates specificity as well as sensitivity: the 1C at $10\,^{\circ}\text{C}$ condition, which exhibited significant capacity fade and whose post-mortem analysis revealed electrolyte decomposition products but no metallic lithium, correctly returned no plating feature, confirming that the GP-identified signature discriminates lithium plating from other degradation pathways. These results demonstrate that the GP–based voltage analysis enables more robust lithium plating diagnostics, paving the way for reliable online detection and integration into advanced battery management systems.
@@ ACKNOWLEDGMENT
We appreciate the funding support from the AFRL Proteus Space STTR and the NSF CAREER Program (No. 2044932).